17. What is the thermal resistance if the composite wall shown has an exposed surface area of 120 m2? a) 9.6 x 105 K/W b) 7.1 x 105 K/W c) 6.3 x 105 K/W d) 2.7 x 10 K/W e 17.9 x 103 K/W -200 62-50 m K m-K 50°C 500°C 200°C 0.5 m
Answer #1

The rate of heat transfer by Fourier's equation is given by:

Q=\frac{kA\Delta T}{L}=\frac{\Delta T}{\frac{L}{kA}}=\frac{\Delta T}{R}

where R is the resistance.

Here the resistance is in series

\therefore R = R_1 +R_2

R_1= \frac{0.3}{200\times120}=1.25\times10^{-5}

R_2= \frac{0.5}{50\times120}=8.33\times10^{-5}

a) R=(1.25+8.33)\times10^{-5}=9.58\times10^{-5} K/W